### v1

# Making the complete OpenAIRE citation graph easily accessible through compact data representation

Abstract
The OpenAIRE graph contains a large citation graph dataset, with over 200 million publications and over 2 billion citations. The current graph is available as a dump with metadata which uncompressed totals $\sim 2.5$ TB. This makes it hard to process on conventional computers. To make this network more available for the community we provide a processed OpenAIRE graph which is downscaled to 32 GB, while preserving the full graph structure. Apart from this we offer the processed data in very simple format, which allows further straightforward manipulation. We also provide a python pipeline, which can be used to process the next releases of the OpenAIRE graph.
Data paper
1 Overview
Repository location
We use two repositories, Zenodo to share our data and fixed version of the processing pipeline, and git repository on Codeberg to facilitate sharing and further development of the processing pipeline.
Data repository
Code repository
Context
The OpenAIRE graph (Rettberg and Schmidt, 2012; Manghi et al., 2025) is a large knowledge graph storing several kinds of research data. In this work we focus on extracting the citations and publications from that dataset, distilling it, and making it more accessible.
There are several other citations graphs; OpenAlex (Priem et al., 2022), Open Research Knowledge Graph (Jaradeh et al., 2019), Crossref (Hendricks et al., 2020) and OpenCitations (Peroni and Shotton, 2020) among others. These datasets differ in many ways, some datasets are only citation graphs, while others like OpenAlex and OpenAIRE are research knowledge graphs that record much more than just citations and publications. The datasets also differ in their publication coverage (Martín-Martín et al., 2021; Culbert et al., 2025) and how the works are indexed, e.g. how they determine taxonomy (Ciuciu-Kiss and Garijo, 2024). Details of publication coverage between OpenAIRE and other open datasets are not well explored and it is not sufficient to compare publication numbers since these datasets aggregate publications from multiple sources, thus a large node number might simply be due to subpar de-duplication.
The OpenAIRE initiative offer several APIs, a search engine and a complete dataset dump (Manghi et al., 2025). The API and search engine serve as accessible avenues for any researcher, however they are limited in the scope of data that can be efficiently accessed. If a researcher is interested in large portions of the graph, or the graph as a whole, then the dump is more suitable.
Processing the whole graph dump, require software coding skills, a large amount of memory, storage and computational power, something not readily accessible to many scholars in the humanities (Dederke et al., 2024). To fill this gap, we pre-process the dump, distilling it into a more managable size, and distributing it as CSVs, making the whole graph, more accessible. Similar work has been done for OpenAlex, to make that citation network more accessible (Caetano Machado Lopes and Chacko, 2024).
2 Method
The OpenAIRE graph consist of many entities and many relations. These are stored in many files in the dump, we only need the publication files (nodes) and the relation files (citations/edges) to extract the citation network. The publication files themselves contain a lot of information about publications, only some of which is relevant for most researchers. The relation files include many types of relations, for this dataset we are only interested in citations (the type ”Cites”). The following shows the generic steps we followed to extract the citation network and minimize memory footprint (the full code is available at [Github link]).
Steps
  1. 1.
    Download the complete OpenAIRE graph dump (Manghi et al., 2025), including all publication and relation files.
  2. 2.
    Extract relevant publication fields and flatten nested JSON structures to obtain a tabular representation.
  3. 3.
    Generate new more memory efficient(int32) nodeIDs, and build a hashtable for translating between OpenAIRE IDs and our nodeIDs.
  4. 4.
    Process relation files using PySpark (Zaharia et al., 2016), retaining only relations of type Cites by extracting source and target identifiers. Use hashtable to replace OpenAIRE IDs
  5. 5.
    Update all citation relations to reference the new identifiers.
This achieves a distilled dataset where the relations are efficiently stored using pairs of short integers.
Quality control
A set of validation scripts are run after processing to check for loss of data. The scripts check whether any entries from the original dump was lost and check for completeness of feature columns.
Output
The processing pipeline takes input of OpenAire full dump and transforms it into the files described below.
  • citations.csv – a simple edge list of the graph (all the citations). Each row has a simple form of two connected nodeIDs of e.g.:
    159486578,118392581
  • publications.csv – all graph nodes (publications), it contains only the nodeID and, when available, the DOI. Each row has a simple form of e.g.:
    14209,10.3931/e-rara-45685
  • publications_large.csv – includes the same number of nodes as publications.csv, but includes additional fields, e.g. title, authors, description, etc. For a full overview of fields, see Table 3.
  • pipeline.zip – processing pipeline which transform full OpenAIRE dump into the dataset files above. Full details is documented in included README.md. The source code is also present online at
    https://codeberg.org/Zmeos/OpenAIRE-citation-extraction.
Memory requirements
Below we summarize the hardware requirements for the transformed dataset and the original OpenAire full dataset dump:
Dataset Size on disk (GB) Memory usage (GB)
citations.csv $39$ $16$
publications.csv 66 $16$
publications_large.csv $179$ $309$
Full OA – edges $1820$ NA
Full OA – nodes $700$ NA
Table 1: Comparison of storage size and memory requirements of the full OpenAIRE (OA) dataset (release 2025-12-01) and its transformed versions. Memory usage corresponds to the amount of GB each dataset occupied when loaded into a pandas dataframe (pandas development team, 2020). Citations are loaded as int32.
3 Dataset Description
Repository name
Zenodo
Object name
An distilled OpenAIRE citation graph
Format names and versions
CSV
Creation dates
Based on the 2025-12-01 OpenAIRE dump.
Dataset creators
Joakim Skarding wrote the pipeline that processed the dump, Pavel Sanda supervised the work. The OpenAIRE initiative, of which this project is unaffiliated, created the OpenAIRE dump.
Language
English
License
CC BY 4.0
Publication date
2026-02-12
4 Reuse Potential
Citation networks are routinely used in a wide range of scientific fields. Among others this includes; research in historical trends of science (Frank et al., 2019; Drivas, 2024; Kitajima and Okamura, 2025; González-Márquez et al., 2024), sociology of scientific knowledge (Crothers et al., 2020; Carradore, 2022), network science (Costa and Frigori, 2024; Xiao et al., 2025) and as training sets for graph neural networks (Kipf, 2016; Leskovec and Sosič, 2016). Since the dataset is an evolving network, it can also be used to train temporal models, such as dynamic graph neural networks (Skarding et al., 2021).
Funding Statement
This work has been funded by a grant from the Programme Johannes Amos Comenius under the Ministry of Education, Youth and Sports of the Czech Republic, CZ.02.01.01/00/23_025/0008711.
Competing interests
The authors have no competing interests to declare.
Supplementary information
Example of reduction of a relation (citation/edge)
Our dump has the benefit of only focusing on citations, thus we do not need to indicate the type of citation. We only retain the information of the source ID and target ID. These IDs are also optimized to take up as little space as possible. The following is an example of how an edge is represented in the OpenAIRE JSON dump:
  "provenance": {
    "provenance": "Inferred by OpenAIRE",
    "trust": "0.9"
  "relType": {
    "name": "Cites",
    "type": "citation"
  "source": "doi_________::7e8d84fc096936557defb78d22cca97c",
  "sourceType": "product",
  "target": "dedup_wf_002::27d83ddfd6e54378d88445aa793d5cb8",
  "targetType": "product",
  "validated": false
An edge in our dump is simply:
159486578,118392581
Estimates of OpenAIRE JSON files uncompressed
Table LABEL:tab:estimate shows a rough estimate of the full uncompressed size of the OpenAIRE dump. The dump would require several orders of magnitude more RAM than most consumer computers.
Table 2: Estimate of OpenAIRE JSON files when uncompressed
Relation (edges) Size Estimate
One relation JSON (uncompressed) 6.5 GB
One relation JSON (compressed) 0.5 GB
Compression ratio $6.5/0.5=13$
Compressed relation folder size 10 GB
Estimated uncompressed per folder $13\times 10=130\,\text{GB}$
Number of relation folders 14
Total uncompressed (relations) $14\times 130=\textbf{1.82\,TB}$
Publication (nodes) Size Estimate
One publication JSON (uncompressed) 0.5 GB
One publication JSON (compressed) 0.1 GB
Compression ratio $0.5/0.1=5$
Compressed publication folder size 10 GB
Estimated uncompressed per folder $5\times 10=50\,\text{GB}$
Number of publication folders 14
Total uncompressed (publications) $14\times 50=\textbf{0.70\,TB}$
Combined Uncompressed Estimate 2.5 TB
Size of datafields in publications_large.csv
Field Explanation Memory usage (GB)
nodeId Unique internal identifier for the node (publication) 22
openaireId Identifier assigned by the OpenAIRE platform $18$
doi Digital Object Identifier of the publication $13$
title Title of the publication $28$
authors List of authors associated with the publication 2020
description Abstract or short description of the publication $192$
date Date when the publication was published $11$
container Journal, conference, or repository where it was published $13$
citations Number of times the publication has been cited 22
language Language in which the publication is written $10$
Table 3: Description of data fields of the large version of the dataset. The publications.csv only contains the nodeId, doi, and citations. The memory usage is the amount of GB that column took when loaded into memory as a pandas dataframe.